Staff

Telescope Operator: Steve allanson

AO support: Doug Miller(LBTO),Juan Carlos Guerra(LBTO)

Astronomer Support: Olga Kuhn

Start

1:41 DIMM:0.76"

WFS Power on: OK

Wind speed:3.5m/s

Temperature: 2C

WFSpower on at 6pm

Summary.

Start Observing.

Problems with the collimation on the left side.GCS is not correcting for high orders. Call Taras and Doug will be reported in the commissioning log.

ACE

Wed Nov 19 02:02:41.296 2014 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[112] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:-0.0673077[-0.0673077],roy:-0.0276644,mag:12.42,cindex:0.87,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

Wed Nov 19 02:03:46.732 2014 aos.PresetAO{FLAO}.complete - PresetAO {FLAO}[112] - dx:nan dy:nan slNull: f1:Filter #0 f2:Filter #0 freq:625 mag:0 nBins:2 nModes:110 r0:0 snMode:0 strehl:0 ttMod:3

Wed Nov 19 02:08:40.816 2014 aos.AcquireRef{FLAO}.failed - AcquireRef {FLAO}[113] - Camera lens centering did not converge after 180 seconds

PRESET FAILED. The System is not receiving frames from the slope computer.

Slope computer and CCD39 reset.

2:35 WFS Power off, w_stop

2:14. DIMM:0.87"

Power on

Wed Nov 19 02:18:30.403 2014 aos.WfsOn.failed - WfsOn [115] - Command timeout - -13107: Arbitrator interface generic error

WfsArbitrator_30101|ERR| 107|2014-11-19 02:19:08.909970| MAIN > Errors for these devices: ['setupXYZstages', 'setupADC']
WfsArbitrator_30101|ERR| 108|2014-11-19 02:19:08.910001| MAIN >
WfsArbitrator_30101|ERR| 109|2014-11-19 02:19:08.910021| MAIN > setupXYZstages: setup did not complete within 120 seconds
WfsArbitrator_30101|ERR| 110|2014-11-19 02:19:08.910040| MAIN > setupADC: Timeout waiting for adc2.L.STATUS

The X-axe didn't home. It is stop at position:78.91, current is -0.02A

Open Eng GUI and home the x-stage from the GUI didn't respond. Diselect/select the Stqage box and the Position value goes to -0.113

Check close loop from Eng GUI. OK.

02:23 close loop test ok

ed Nov 19 02:24:26.208 2014 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[117] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:-0.0408882[-0.0408882],roy:-0.0305155,mag:12.42,cindex:0.87,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

Looking hte WFS Eng GUI the bayside stage-X is not moving.

Wed Nov 19 02:26:36.272 2014 aos.PresetAO{FLAO}.failed - PresetAO {FLAO}[117] - Command timeout - -13107: Arbitrator interface generic error

rbitrator_30101|ERR| 207|2014-11-19 02:27:18.449044| MAIN > Timeout error on baysidex (-5001) TIMEOUT_ERROR
WfsArbitrator_30101|ERR| 208|2014-11-19 02:27:18.449094| MAIN > Traceback (most recent call last):
WfsArbitrator_30101|ERR| 209|2014-11-19 02:27:18.449115| MAIN > File "/home/aoacct/AO/current/lib/python/AdOpt/WfsArbScripts.py", line 2508, in arbFunc
WfsArbitrator_30101|ERR| 210|2014-11-19 02:27:18.449134| MAIN > errstr = func(*args)
WfsArbitrator_30101|ERR| 211|2014-11-19 02:27:18.449153| MAIN > File "/home/aoacct/AO/current/lib/python/AdOpt/WfsArbScripts.py", line 919, in doPrepareAcquireRef
WfsArbitrator_30101|ERR| 212|2014-11-19 02:27:18.449171| MAIN > app.stagex.waitTargetReached()
WfsArbitrator_30101|ERR| 213|2014-11-19 02:27:18.449189| MAIN > File "/home/aoacct/AO/current/lib/python/AdOpt/hwctrl/simplemotor.py", line 300, in waitTargetReached
WfsArbitrator_30101|ERR| 214|2014-11-19 02:27:18.449209| MAIN > raise AdOptError ( AOErrCode ['TIMEOUT_ERROR'], "Timeout error on " Kelf.name)
WfsArbitrator_30101|ERR| 215|2014-11-19 02:27:18.449227| MAIN > AdOptError: Timeout error on baysidex - code: -5001 TIMEOUT_ERROR

Move manually OK.

Wed Nov 19 02:28:07.184 2014 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[119] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:-0.0364576[-0.0364576],roy:-0.030858,mag:12.42,cindex:0.87,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

Wed Nov 19 02:33:32.817 2014 aos.info - WFS: star found, estimated magnitude: 13.76Wed Nov 19 02:34:24.423 2014 aos.AcquireRef{FLAO}.complete - AcquireRef {FLAO}[120] - dx:0 dy:0 slNull: f1:Filter #0 f2:Filter #0 freq:200 mag:13.76 nBins:3 nModes:66 r0:0 snMode:0 strehl:0 ttMod:6

Wed Nov 19 02:34:24.445 2014 aos.StartAO{FLAO}.started - StartAO {FLAO}[121]

The image in the CCD47 was moving in circle. Stop the ADC

ADC power on and the image in the CCD47 stop rotating.

TN:#20141119_023927,_024056,_024125,_024328,_025906

2:41 DIMM?
Luci set:20141119_0005-0010

Wed Nov 19 03:03:10.417 2014 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[159] (aomode:TTM-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0.00545615[0.00545615],roy:-0.0321179,mag:12.42,cindex:0.87,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

Wed Nov 19 03:03:58.518 2014 aos.PresetAO{FLAO}.complete - PresetAO {FLAO}[159] - dx:nan dy:nan slNull: f1:Filter #0 f2:Filter #0 freq:400 mag:0 nBins:4 nModes:10 r0:0 snMode:0 strehl:0 ttMod:6

image bad in the ccd47, something wrong with the offloading of high orders to teh primary. GCS is not sending corrections.

TN:_032255,_032312

3:25 Collimation issues on the left side. Doug is working with the IDL to do the collimation.

TTM

Wed Nov 19 03:39:18.125 2014 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[165] (aomode:TTM-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0.0488256[0.0488256],roy:-0.0298354,mag:12.42,cindex:0.87,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

Wed Nov 19 03:39:51.190 2014 aos.PresetAO{FLAO}.complete - PresetAO {FLAO}[165] - dx:nan dy:nan slNull: f1:Filter #0 f2:Filter #0 freq:400 mag:0 nBins:4 nModes:10 r0:0 snMode:0 strehl:0 ttMod:6

Wed Nov 19 03:46:14.967 2014 aos.StartAO{FLAO}.complete - StartAO {FLAO}[167]

TN:#_034642,_034708,_034748,_03525
Luci set:20141119_0011-0016

3:54.DIMM:0.85"

Active

4:01 DIMM:0.82"
GCS set: _234-
Luci set:20141119_0017-0022
CCD47 test.

TN:#_041736 (10 frames)

ACE

Wed Nov 19 04:19:01.499 2014 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[206] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0.0282448[0.0282448],roy:-0.167451,mag:15,cindex:-4.98,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

Wed Nov 19 04:20:15.781 2014 aos.PresetAO{FLAO}.complete - PresetAO {FLAO}[206] - dx:nan dy:nan slNull: f1:Filter #0 f2:Filter #0 freq:200 mag:0 nBins:4 nModes:36 r0:0 snMode:0 strehl:0 ttMod:6

No Data collected.

TTM

Wed Nov 19 04:30:14.298 2014 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[208] (aomode:TTM-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:-29.9155[-29.9155],roy:15.8959,mag:14.21,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

Wed Nov 19 04:33:08.456 2014 aos.info - WFS: star found, estimated magnitude: 14.82

Wed Nov 19 04:33:36.046 2014 aos.info - Adsec 672b-left loading reconstructor.

Wed Nov 19 04:33:37.716 2014 aos.info - Stages offset applied - dx: 0; dy: 0

Wed Nov 19 04:33:37.742 2014 aos.AcquireRef{FLAO}.complete - AcquireRef {FLAO}[209] - dx:0 dy:0 slNull: f1:Filter #0 f2:Filter #0 freq:200 mag:14.824 nBins:4 nModes:10 r0:0 snMode:0 strehl:0 ttMod:6

Wed Nov 19 04:33:37.791 2014 aos.StartAO{FLAO}.started - StartAO {FLAO}[210]

TN:#_043606,_043621,_043739,_043753
Luci set:0025-0030

TN:#_045032,_045128

Active

Luci set:0031-0036
CCD47 test.

TN:#045920,_053205,_053224,_
WFS set: ?

5:5 DIMM:0.8"

TTM

Wed Nov 19 05:35:37.645 2014 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[260] (aomode:TTM-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0.143244[0.143244],roy:0.168194,mag:11.86,cindex:0.89,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

DAta taken hold.

05:42. Some issue in the right side MODS2 guider. Working on the right side..

Error taking ccd47 images. The system is AOPrepared, Adsec: AOSET. The ccd47 is changed to bin 1

WfsArbitrator _30101|DEB| 1711|2014-11-19 05:41:33.268858| MAIN > arbFunc savePsfInternal (, 'FLAOWFS', 'LUCIFER', '/towerdata/adsec_data/20141119/Data_20141119_054132/psf.fits', 10)
WfsArbitrator_30101|INF| 1712|2014-11-19 05:41:33.268897| MAIN > savePsfInternal(): starting measure, 10 frames
WfsArbitrator_30101|ERR| 1713|2014-11-19 05:41:33.857628| MAIN > RETRY: Cannot save PSF (0) NO_ERROR
WfsArbitrator_30101|ERR| 1714|2014-11-19 05:41:33.857661| MAIN > Traceback (most recent call last):
WfsArbitrator_30101|ERR| 1715|2014-11-19 05:41:33.857685| MAIN > File "/home/aoacct/AO/current/lib/python/AdOpt/WfsArbScripts.py", line 2508, in arbFunc
WfsArbitrator_30101|ERR| 1716|2014-11-19 05:41:33.857705| MAIN > errstr = func(*args)

05:48 Off-axis guiding problem on the right side. John H called.

Change to bin 2 and we have the same error. The bin16 is working and save data from the optical loop data gui

-- %USERSIG{JuanCarlosGuerra - 2014-11-19}%

6:02. HAND OVER TO LBTI

9:52 BACK

Open loop images, telescope not tracking
Luci Set: 0039

temperature :2

wind speed:3m/s

some cirrus passing

Check pointing and collimation

TTM

Wed Nov 19 10:08:55.676 2014 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[417] (aomode:TTM-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:3.61806[3.61806],roy:-5.01729,mag:11.04,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

Wed Nov 19 10:09:34.104 2014 aos.PresetAO{FLAO}.complete - PresetAO {FLAO}[417] - dx:nan dy:nan slNull: f1:Filter #0 f2:Filter #0 freq:400 mag:0 nBins:4 nModes:10 r0:0 snMode:0 strehl:0 ttMod:6

10:22 DIMM:0.74"

Taking images for Luci filter H,H2,Br_gm and Ks,J
Luci set:0041-0050

Wed Nov 19 10:47:21.449 2014 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[432] (aomode:TTM-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:3.66767[3.66767],roy:-5.00671,mag:11.04,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

Wed Nov 19 10:47:54.142 2014 aos.PresetAO{FLAO}.complete - PresetAO {FLAO}[432] - dx:nan dy:nan slNull: f1:Filter #0 f2:Filter #0 freq:400 mag:0 nBins:4 nModes:10 r0:0 snMode:0 strehl:0 ttMod:6

Preset Cancel.

The problem is the right side.

Wed Nov 19 10:57:23.635 2014 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[437] (aomode:TTM-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:3.68133[3.68133],roy:-5.00303,mag:11.04,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Luci set:0051-0062

offset 120" sky:

Stop Preset, problems with the primary on the right side.

Wed Nov 19 11:37:40.406 2014 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[520] (aomode:TTM-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:3.74089[3.74089],roy:-4.98354,mag:11.04,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

stop.

New preset.

Active

Sky images.

Luci Set: 0064-0067

Change filter to H2
Luci Set: 0068-0072

TTM

Wed Nov 19 12:11:05.255 2014 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[528] (aomode:TTM-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:3.79618[3.79618],roy:-4.9605,mag:11.04,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

TN:#_121536,_121554,_121643,_121755
Luci Set: 0073-0077

Change filter Br_g
Luci Set: 0078-0082,0083-87

12:32.DIMM:0.78"

12:48 DIMM:0.9"

tn:#125026,_125114,

Go to Sky

ACTIVE

Luci Set: 0088-0092

NIGHT NIGHT 13:17

darks Luci:4sec: 0093,0094,0095,0096

6sec:0097,0098,0099,0100,0101

4sec0102-0106

Comments

 
Topic revision: r3 - 19 Nov 2014, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback